# IAP: Improving Continual Learning of Vision-Language Models via Instance-Aware Prompting

### v1

@@ Abstract
Recent pre-trained vision-language models (PT-VLMs) often face a Multi-Domain Class-Incremental Learning (MCIL) scenario in practice, where several classes and domains of multi-modal tasks are incrementally arrived. Without access to previously learned tasks and unseen tasks, memory-constrained MCIL suffers from forward and backward forgetting. To alleviate the above challenges, parameter-efficient fine-tuning techniques (PEFT), such as prompt tuning, are employed to adapt the PT-VLM to the diverse incrementally learned tasks. To achieve effective new task adaptation, existing methods only consider the effect of PEFT strategy selection, but neglect the influence of PEFT parameter setting (e.g., prompting). In this paper, we tackle the challenge of optimizing prompt designs for diverse tasks in MCIL and propose an Instance-Aware Prompting (IAP) framework. Specifically, our Instance-Aware Gated Prompting (IA-GP) module enhances adaptation to new tasks while mitigating forgetting by dynamically assigning prompts across transformer layers at the instance level. Our Instance-Aware Class-Distribution-Driven Prompting (IA-CDDP) improves the task adaptation process by determining an accurate task-label-related confidence score for each instance. Experimental evaluations across 11 datasets, using three performance metrics, demonstrate the effectiveness of our proposed method. Code can be found at https://github.com/FerdinandZJU/IAP.
1 Introduction
@@
Recent years have witnessed a great development of deep neural networks in numerous multi-modal applications, where all the required data are available simultaneously for training on various tasks [52]. Nevertheless, real-world applications usually meet a challenging Multi-Domain Class-Incremental Learning (MCIL) scenario [48]: 1) tasks of multi-modal data are acquired incrementally, 2) the task distribution and target classes vary across tasks. Since the learned tasks data are un-available in a memory-constrained MCIL, multi-modal models suffer from the catastrophic forgetting phenomenon [52] from two aspects: backward forgetting and forward forgetting [51], as shown in Figure 1.
Figure 1: Illustration of backward forgetting and forward forgetting. During the current learning process (e.g., Flowers), backward forgetting in traditional incremental learning refers to the degradation of seen tasks (e.g., Aircraft and DTD). In contrast, forward forgetting manifests as a decline in zero-shot generalization performance on unseen tasks (e.g., Food and SUN397).
@@
Due to the remarkable zero-shot ability of vision-language models (VLMs) [12] (e.g., CLIP [33]), existing works have explored the multi-modal class incremental learning with pre-trained VLMs and parameter-efficient fine-tuning (PEFT) techniques (e.g., prompt, adapter, LoRA) [52]. Typically, these pre-trained model-based multi-modal class incremental learning works usually maintain a fixed pre-trained VLM and incrementally learn few task-specific parameters to adapt to various encountered tasks. Since the memory of MCIL is constrained, such pre-trained model-based methods mainly focus on designing an appropriate PEFT strategy for effective adaptation on various tasks. Both the type of PEFT strategy (prompt, adapter or LoRA) and the parameter setting of PEFT (where and how many prompts/adapters/LoRA to integrate) influence the performance of learning a new task in MCIL.
@@
Existing MCIL works have widely discussed the effect of PEFT strategy selection on MCIL performance, but overlook the analysis of the PEFT parameter settings. These works typically employ a manually designed PEFT parameter setting and apply them consistently across the incremental learning of different tasks. For instance, when learning a new task, prompt-based MCIL methods (e.g., DIKI [36]) consistently use the same number of prompts at the same positions within the pre-trained VLM (an operation we refer to as prompting). However, our observations suggest that the optimal prompting strategy for adapting to different tasks varies within the MCIL framework. Thus, the key challenge lies in developing a flexible and adaptive prompting strategy to enhance task adaptation.
@@
Motivated by the above observations, we propose a novel Instance-Aware Prompting (IAP) framework for MCIL, which dynamically allocates prompt positions and weights at the instance level. Specifically, we introduce an Instance-Aware Gated Prompting (IA-GP) strategy that determines whether to apply prompts in each transformer layer based on instance-specific features. Furthermore, we propose an Instance-Aware Class-Distribution-Driven Prompting (IA-CDDP) strategy to compute more reliable confidence scores at the instance level, which are then used as prompt weights in subsequent operations. By integrating above two modules, we solve the problems of fixed prompting strategy of conducting incremental learning on pre-trained VLMs.
@@
  • We introduce Instance-Aware Class-Distribution-Driven Prompting to derive more reliable prompt weights. To enhance model performance across diverse instances, we employ a two-stage distribution modeling approach that operates at both the task and class levels during inference.
@@
  • Extensive experiments on benchmark datasets demonstrate that our IAP method achieves state-of-the-art performance while utilizing only 1.18% of the training parameters, outperforming all existing approaches.
2 Related Works
@@
Incremental learning. Conventional incremental learning can be categorized into three types: task-incremental learning (TIL) [17, 38, 51], domain-incremental learning (DIL) [17, 38, 50], and class-incremental learning (CIL) [17, 38, 49, 11]. Among these, task-incremental learning requires explicit task identifies during inference, while class-incremental learning requires non-overlapping classes across different tasks. There are three technologies in incremental learning. The first is regularization-based methods, such as Elastic Weight Consolidation (EWC) [20] and Memory Aware Synapses (MAS) [1]. These methods introduce regularization terms to constrain optimization directions and mitigate catastrophic forgetting. The second one is rehearsal-based methods [26, 34, 43, 15, 24, 16, 30], including Learning without Forgetting (LwF) [26] and Incremental Classifier and Representation Learning (iCaRL) [34]. These approaches generally require additional storage to retain either previous model parameters or representative samples from seen tasks. The third one is network expansion-based methods [29, 46, 45, 25], with Dynamic Expandable Networks (DEN) [46] being a representative approach. These methods dynamically expand the neural network structure to accommodate new tasks. Recently, observing the strong generalization of pre-trained models, some incremental learning approaches are designed to fine-tune the pre-trained models [19, 37, 53, 4, 13], which called parameter-efficient fine-tuning (PEFT). PEFT techniques achieve adaptation for downstream tasks by fine-tuning only a small number of model parameters or maintaining a limited set of additional parameters while keeping the majority of the pre-trained model’s parameters frozen. PEFT technology significantly reduces computational and storage costs and mitigates the catastrophic forgetting. Notable related methods include L2P [41], DualPrompt [40], S-Prompt [39], and CODA-Prompt [35]. However, these methods only consider stable distributions and lack of the capability to learn tasks with distinct distributions. Unlike conventional approaches, we focus on a novel benchmark called Multi-Domain Class-Incremental Learning (MCIL), where data distributions are distinct and arrive continually.
@@
Downstream tasks of vision-language models. Visual Language Models (VLMs) [18, 32] have achieved remarkable progress in multimodal research, successfully enabling cross-modal understanding and generation capabilities through joint modeling of visual inputs and natural language. With the explosive growth of online data, recent years have witnessed further advancements in VLMs driven by larger-scale models and more extensive datasets. However, when researchers attempt to fine-tune the VLMs for a downstream task, an inevitable degradation occurs in their zero-shot capabilities of other tasks [23, 5]. Various approaches have been proposed to mitigate the degradation. ZSCL [51] leverages the knowledge distillation technology [26], treating the original VLM as a teacher model and distilling knowledge into the fine-tuned model through constructed “wild” [3] datasets (e.g., ImageNet [7]). Although ZSCL partially alleviates forgetting of pre-trained knowledge, it requires additional storage space for “wild” datasets, and its full-parameter fine-tuning strategy incurs substantial computational costs. Alternative approaches such as MoE-Adapter [47] and DIKI [36] employ PEFT technology, updating only a small set of parameters to reduce the forgetting of pre-trained knowledge. Nevertheless, these methods continue to encounter challenges associated with computational resource overhead and the under-exploration of instance-aware features. Our proposed approach overcomes these limitations by introducing an instance-aware prompting strategy, significantly enhancing model performance.
Figure 2: Illustration of Instance-Aware Gated Prompting strategy. The figure illustrates the processing workflow of the IA-GP strategy applied to an instance ”abbey”. Initially, the image is segmented into patches, which are then fed into the visual encoder of the CLIP model. Simultaneously, corresponding textual category descriptions are processed by the textual encoder. The IA-GP strategy utilizes trainable prompt pools for both visual and textual modalities, while prompts from previously seen tasks are kept frozen. The strategy leverages the visual features from the original CLIP encoders, denoted as $f_{v}=F_{v}(x)$, as input of prompt gate module. Hard Gumbel logits produced by prompt gate modules are used to determine whether to retrieve prompts from prompt pools. $L$ denoted each self-attention layer in the Transformer architectures. Outputs from the original CLIP model are represented in blue, while those incorporating retrieved prompts via the IA-GP strategy are shown in green. The processed visual and textual features are optimized through a contrastive loss function.
3 Approach
3.1 Preliminaries
@@
Mutli-Domain Class-Incremental Learning. Given a pre-trained VLM, which incrementally learn from a stream of tasks originating from $\mathcal{T}$ distinct domains ${\{D_{1},D_{2},...,D_{\mathcal{T}}\}}$, each domain consists of ${N}$ samples, denoted as $D_{t}=(x^{t}_{n},y^{t}_{n})^{N}_{n=1}$, where ${x_{n}}$ is an image and $y_{n}$ is the corresponding one-hot ground truth. For domain $D_{n}$, the associated class set is given by $C_{n}=\{c^{n}_{i}\}^{M^{n}}_{i=1}$, where each $c^{n}_{i}$ represents a textual label. Under MCIL setting, the domain ${D_{n}}$ is accessible only during its corresponding $n$-th incremental learning phase. Moreover, class sets are disjoint across domains, i.e., $C_{i}\neq C_{j}$ for any $i\neq j$, and the data distributions also differ across domains, meaning $\mathbb{P}_{i}\neq\mathbb{P}_{j}$, where $\mathbb{P}$ denotes the distribution. During the test phase, the model is evaluated without accessing to the task identifier, requiring it to infer predictions across all previously seen domains without explicit domain information.
@@
CLIP model. The pre-trained VLMs such as CLIP [33] consisted of an image encoder $F_{v}$ and a text encoder $F_{t}$. For a class $c_{i}^{n}$, CLIP model first transforms it to a sentence by a template such as “{a photo of {$c_{i}^{n}$}}”, and then encodes it into text embedding $t_{i}$. CLIP model is trained by leverages contrastive loss, the objective can be defined as:
@@
Interference-free Knowledge Integration (IKI) mechanism. We follow  [36] to leverage a prompt-based incremental learning method. Specifically, a set of prompt pools are maintained for a stream of tasks, which can be denoted as:
$\mathcal{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},\dots,\mathcal{P}_{\mathcal{T}}\},$ (2)
where $P_{t}=(K_{t},V_{t})$, and $K_{t},V_{t}\in\mathbb{R}^{l\times d}$. $l$ is the prompt length, $d$ is the embedding dimension. When a test sample $x_{n}$ comes, IKI mechanism first selects the corresponding prompt $\mathcal{P}_{r}$, and produces a task-specific attention output, which can be formulated as:
@@
3.2 Instance-Aware Prompting
3.2.1 Gated Prompting
@@
The IA-GP strategy incorporates multiple prompting gates positioned before the vision transformer layers, as illustrated in Figure 2. Each prompt gate consists of a Gumbel linear function $H_{i}$, which maps the embedded image features to a $K$-dimensional space, and a Gumbel distribution which used to generate the samples uniformly. For each instance, we utilize the image features extracted by the original CLIP model as the input to the IA-GP module. This operation is motivated by the fact that the original CLIP model is frozen, ensuring that its raw image features reside in a stable distribution. We then compute the instance-aware Gumbel logit by:
@@
In our approach, we set $K=2$ to implement a hard Gumbel Softmax mechanism, which facilitates binary decision-making for prompting. $g_{i}=-\log(-\log(U_{i}))$ represents a random logit sampled from the Gumbel distribution, where $U_{i}\sim U(0,1)$. The parameter $\tau$ denotes the temperature of the IA-GP module, controlling the sharpness of the decision boundary. The IA-GP module generates a two-dimensional output $G_{i}$, which is subsequently transformed into a one-hot vector via the hard Gumbel Softmax operation. This mechanism enables IAP approach to dynamically determine whether to prompt for each layer at the instance level. Specifically, if prompting is beneficial, the IA-GP module adjusts the Gumbel logit such that $G_{i}$ approaches 1. Conversely, if prompting operation is detrimental, the gate sets $G_{i}$ close to 0, thereby preserving the original output of the CLIP model. In summary, for each transformer layer during training, the output under our IA-GP strategy is formulated as follows:
@@
The operations described above are implemented for each individual instance. As a result, the IA-GP module can dynamically determining and assigning the appropriate the prompting layers according to the specific features of each instance.
3.2.2 Class-Distribution Driven Prompting
Incremental VLMs must mitigate backward forgetting (degradation of learned distributions) and prevent forward forgetting (impaired generalization to unseen distributions) to preserve zero-shot generalization. In the MCIL scenario, where task identifiers are absent during inference, distinguishing seen tasks from unseen tasks is crucial. DIKI [36] adjusts prompt weights using instance-to-task similarity as a confidence score. However, this approach is suboptimal, where high-confidence instances should use unadjusted prompts directly, while low-confidence instances risk noise if prompted, the original CLIP model needs to be employed. For instances with intermediate confidence scores, the confidence score needs to be more reliable. To address this, we propose Instance-Aware Class-Distribution-Driven Prompting (IA-CDDP), a two-stage strategy that reassesses instances via both task- and class-specific perspectives during inference.

### v2

@@ Abstract
Recent pre-trained vision-language models (PT-VLMs) often face a Multi-Domain Task Incremental Learning (MTIL) scenario in practice, where several classes and domains of multi-modal tasks are incrementally arrived. Without access to previously seen tasks and unseen tasks, memory-constrained MTIL suffers from forward and backward forgetting. To alleviate the above challenges, parameter-efficient fine-tuning techniques (PEFT), such as prompt tuning, are employed to adapt the PT-VLM to the diverse incrementally learned tasks. To achieve effective new task adaptation, existing methods only consider the effect of PEFT strategy selection, but neglect the influence of PEFT parameter setting (e.g., prompting). In this paper, we tackle the challenge of optimizing prompt designs for diverse tasks in MTIL and propose an Instance-Aware Prompting (IAP) framework. Specifically, our Instance-Aware Gated Prompting (IA-GP) strategy enhances adaptation to new tasks while mitigating forgetting by adaptively assigning prompts across transformer layers at the instance level. Our Instance-Aware Class-Distribution-Driven Prompting (IA-CDDP) improves the task adaptation process by determining an accurate task-label-related confidence score for each instance. Experimental evaluations across 11 datasets, using three performance metrics, demonstrate the effectiveness of our proposed method. The source codes are available at https://github.com/FerdinandZJU/IAP.
Index Terms:
Continual learning, Incremental learning, Multi-modal learning
I Introduction
@@
Recent years have witnessed a great development of deep neural networks in numerous multi-modal applications, where all the required data are available simultaneously for training on various tasks [1, 2]. Nevertheless, real-world applications usually meet a challenging Multi-Domain Task Incremental Learning (MTIL) scenario [3]: 1) tasks of multi-modal data are acquired incrementally, 2) the task distribution and target classes vary across tasks. Since the learned tasks data are un-available in a memory-constrained MTIL, multi-modal models suffer from the catastrophic forgetting phenomenon [2] from two aspects: backward forgetting and forward forgetting [3], as shown in Figure 1.
@@
Due to the remarkable zero-shot ability of vision-language models (VLMs) [4, 5, 6, 7] (e.g., CLIP [8]), existing works have explored the multi-modal class incremental learning with pre-trained VLMs and parameter-efficient fine-tuning (PEFT) techniques (e.g., prompt, adapter, LoRA) [2, 1]. Typically, these pre-trained model-based multi-modal class incremental learning works usually maintain a fixed pre-trained VLM and incrementally learn few task-specific parameters to adapt to various encountered tasks. Since the memory of MTIL is constrained, such pre-trained model-based methods mainly focus on designing an appropriate PEFT strategy for effective adaptation on various tasks. The selection of PEFT strategy—whether prompt, adapter, or LoRA—as well as the configuration of PEFT parameters, including the placement and quantity of prompts, adapters, or LoRA modules, significantly influences the performance of MTIL for acquiring new tasks.
@@
Existing MTIL works have widely discussed the effect of PEFT strategy selection on MTIL performance, but overlook the analysis of the PEFT parameter settings. These works typically employ a manually designed PEFT parameter setting and apply them consistently across the incremental learning of different tasks. For instance, when learning a new task, prompt-based MTIL methods (e.g., DIKI [9]) consistently use the same number of prompts at the same positions within the pre-trained VLM (an operation we refer to as prompting). However, our observations suggest that the optimal prompting strategy for adapting to different tasks varies within the MTIL framework. Thus, the key challenge lies in developing a flexible and adaptive prompting strategy to enhance task adaptation.
Figure 1: Illustration of backward forgetting and forward forgetting. During the current learning process (e.g., Flowers), backward forgetting refers to the degradation of seen tasks (e.g., Aircraft and DTD); forward forgetting refers to the decline in zero-shot generalization capability on unseen tasks (e.g., Food and SUN397).
@@
Motivated by the above observations, we propose a novel Instance-Aware Prompting (IAP) framework for MTIL, which adaptively allocates prompt positions and weights at the instance level. Specifically, we introduce an Instance-Aware Gated Prompting (IA-GP) strategy that determines whether to apply prompts in each transformer layer based on instance-specific features. Furthermore, we propose an Instance-Aware Class-Distribution-Driven Prompting (IA-CDDP) strategy to compute more reliable confidence scores at the instance level, which are then used as prompt weights in subsequent operations. By integrating above two strategies, we solve the problems of fixed prompting strategy of conducting incremental learning on pre-trained VLMs.
@@
  • We introduce an Instance-Aware Class-Distribution-Driven Prompting strategy to derive more reliable prompt weights. To enhance model performance across diverse instances, we employ a two-stage distribution modeling strategy that operates at both the task and class levels during inference.
@@
  • Extensive experiments on benchmark datasets demonstrate that our method achieves state-of-the-art (SOTA) performance while utilizing only 1.18% of the training parameters, outperforming all existing methods.
II Related Works
II-A Incremental Learning
@@
Conventional incremental learning can be categorized into three types: task-incremental learning (TIL) [10, 11, 3, 12], domain-incremental learning (DIL) [13], and class-incremental learning (CIL) [14, 15, 16, 17, 18, 19, 20]. Among these, task-incremental learning requires explicit task identifies during inference, while class-incremental learning requires non-overlapping classes across different tasks. Three technologies are generally utilized in incremental learning: 1) Regularization-based methods, such as Elastic Weight Consolidation (EWC) [21] and Memory Aware Synapses (MAS) [22]. These methods introduce regularization terms to constrain optimization directions and mitigate catastrophic forgetting. 2) Rehearsal-based methods [23, 24, 25, 26, 27], Representative methods are Learning without Forgetting (LwF) [28] and Incremental Classifier and Representation Learning (iCaRL) [29]. These approaches require additional storage to retain either previous model parameters or representative samples from seen tasks. 3) Network expansion-based methods [30, 31], with Dynamic Expandable Networks (DEN) [32] being a representative approach. These methods dynamically expand the neural network structure to accommodate new tasks. Recently, observing the strong generalization of pre-trained models, some incremental learning approaches are designed to fine-tune the pre-trained models [33, 34, 35, 36, 37], which called parameter-efficient fine-tuning (PEFT). PEFT techniques achieve adaptation for downstream tasks by fine-tuning only a small number of model parameters or maintaining a limited set of additional parameters while keeping the majority of the pre-trained model’s parameters frozen. PEFT technology significantly reduces computational and storage costs and mitigates the catastrophic forgetting. Notable related methods include L2P [38], DualPrompt [39], S-Prompt [40], and CODA-Prompt [41]. However, these methods only consider stable distributions and lack of the capability to learn tasks with distinct distributions. Our work focus on MTIL benchmark, where data distributions are distinct and arrive continually.
II-B Downstream tasks of vision-language models
@@
Visual Language Models (VLMs) [42, 43] have achieved remarkable progress in multi-modal research, successfully enabling cross-modal understanding and generation capabilities through joint modeling of visual inputs and natural language. With the explosive growth of online data, recent years have witnessed further advancements in VLMs driven by larger-scale models and more extensive datasets. However, when researchers attempt to fine-tune the VLMs for a downstream task, an inevitable degradation occurs in their zero-shot capabilities of other tasks [44, 45]. Various approaches have been proposed to mitigate the degradation. ZSCL [3] leverages the knowledge distillation technology [28], treating the original VLM as a teacher model and distilling knowledge into the fine-tuned model through constructed “wild” [46] datasets (e.g., ImageNet [47]). Although ZSCL partially alleviates forgetting of pre-trained knowledge, it requires additional storage space for “wild” datasets, and its full-parameter fine-tuning strategy incurs substantial computational costs. Alternative approaches such as MoE-Adapter [48] and DIKI [9] employ PEFT technology, updating only a small set of parameters to reduce the forgetting of pre-trained knowledge. Nevertheless, these methods continue to encounter challenges associated with computational resource overhead and the under-exploration of instance-aware features. Our proposed approach overcomes these limitations by introducing an instance-aware prompting strategy, significantly enhancing model performance.
III Methodology
III-A Preliminaries
III-A1 MTIL Benchmark
@@
Given a pre-trained VLM, which incrementally learns from a stream of tasks originating from $\mathcal{T}$ distinct domains ${\{D_{1},D_{2},...,D_{\mathcal{T}}\}}$, each domain consists of ${N}$ samples, denoted as $D_{t}=(x^{t}_{n},y^{t}_{n})^{N}_{n=1}$, where ${x_{n}}$ is an image and $y_{n}$ is the corresponding one-hot ground truth. For domain $D_{n}$, the associated class set is given by $C_{n}=\{c^{n}_{i}\}^{M^{n}}_{i=1}$, where each $c^{n}_{i}$ represents a textual label. In the MTIL setting, the domain ${D_{n}}$ is accessible only during its corresponding $n$-th incremental learning session. Moreover, class sets are disjoint across domains, i.e., $C_{i}\neq C_{j}$ for any $i\neq j$, and the data distributions also distinct across domains, meaning $\mathbb{P}_{i}\neq\mathbb{P}_{j}$, where $\mathbb{P}$ denotes the distribution. During the inference, the model is evaluated without accessing to the task identifier, requiring it to infer across all previously seen domains without explicit domain information.
III-A2 CLIP Model
@@
The pre-trained VLMs such as CLIP [8] consisted of an image encoder $F_{v}$ and a text encoder $F_{t}$. For a class $c_{i}^{n}$, CLIP model first transforms it to a sentence by a template such as “{a photo of {$c_{i}^{n}$}}”, and then encodes it into text embedding $t_{i}$. CLIP model is trained by leverages contrastive loss, the objective can be defined as:
@@
III-A3 Interference-Free Knowledge Integration (IKI) Mechanism
We follow  [9] to leverage a prompt-based incremental learning method. Specifically, a set of prompt pools are maintained for a stream of tasks, which can be denoted as $\mathcal{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},\dots,\mathcal{P}_{\mathcal{T}}\}$, where $P_{t}=(K_{t},V_{t})$, and $K_{t},V_{t}\in\mathbb{R}^{l\times d}$. $l$ is the prompt length, $d$ is the embedding dimension. When a test sample $x_{n}$ comes, IKI mechanism first selects the corresponding prompt $\mathcal{P}_{r}$, and produces a task-specific attention output, which can be formulated as:
@@ III-A3 Interference-Free Knowledge Integration (IKI) Mechanism
Figure 2: Illustration of Instance-Aware Gated Prompting strategy. The figure illustrates the processing workflow of the IA-GP strategy applied to an instance “abbey”. The image is segmented into patches initially, which are then fed into the visual encoder of the CLIP model. The prompt pools of seen tasks are kept frozen. IA-GP leverages the visual features from the original CLIP encoders, denoted as $f_{v}=F_{v}(x)$, as input of prompt gate module. Hard Gumbel logits produced by prompt gate modules are used to determine whether to retrieve prompts from prompt pools. $L$ denoted each self-attention layer in the Transformer architectures. Outputs from the original CLIP model are represented in blue, while those incorporating retrieved prompts via the IA-GP strategy are shown in green. The processed visual and textual features are optimized through a contrastive loss function.
III-B Instance-Aware Prompting
III-B1 Gated Prompting
@@
The IA-GP strategy incorporates multiple prompting gates positioned before the vision transformer layers, as illustrated in Figure 2. Each prompt gate consists of a Gumbel linear function $H_{i}$, which maps the embedded image features to a $K$-dimensional space, and a Gumbel distribution which used to generate the samples uniformly. For each instance, we utilize the image features extracted by the original CLIP model as the input to the IA-GP strategy. This operation is motivated by the fact that the original CLIP model is frozen, ensuring that its raw image features reside in a stable distribution. We then compute the instance-aware Gumbel logit by:
@@
In our approach, we set $K=2$ to implement a hard Gumbel Softmax mechanism, which facilitates binary decision-making for prompting. $g_{i}=-\log(-\log(U_{i}))$ represents a random logit sampled from the Gumbel distribution, where $U_{i}\sim U(0,1)$. The parameter $\tau$ denotes the temperature of the IA-GP strategy, controlling the sharpness of the decision boundary. The IA-GP strategy generates a two-dimensional output $G_{i}$, which is subsequently transformed into a one-hot vector via the hard Gumbel Softmax operation. This mechanism enables IAP approach to dynamically determine whether to prompt for each layer at the instance level. Specifically, if prompting is beneficial, the IA-GP strategy adjusts the Gumbel logit such that $G_{i}$ approaches 1. Conversely, if prompting operation is detrimental, the gate sets $G_{i}$ close to 0, thereby preserving the original output of the CLIP model. In summary, for each transformer layer during training, the output of our IA-GP strategy is formulated as follows:
@@
The operations described above are implemented for each individual instance. As a result, the IA-GP strategy can dynamically determine and assign the appropriate prompting layers according to the specific features of each instance.
III-B2 Class-Distribution Driven Prompting
Incremental VLMs must mitigate backward forgetting (degradation of learned distributions) and prevent forward forgetting (impaired generalization to unseen distributions) to preserve zero-shot generalization. In the MTIL scenario, where task identifiers are absent during inference, distinguishing seen tasks from unseen tasks is crucial. DIKI [9] adjusts prompt weights using instance-to-task similarity as a confidence score. However, this approach is suboptimal, where high-confidence instances should use unadjusted prompts directly, while low-confidence instances risk noise if prompted, the original CLIP model needs to be employed. For instances with intermediate confidence scores, the confidence score needs to be more reliable. To address this, we propose Instance-Aware Class-Distribution-Driven Prompting (IA-CDDP), a two-stage strategy that reassesses instances via both task- and class-specific perspectives during inference.